a. A transformer is used to change the 230 volt mains electricity supply to the 12 volts

needed to operate a low voltage halogen lamp. The current through the halogen lamp

is 4 amps. Assuming that the transformer is 100 % efficient determine the:

ii. number of turns on the primary coil if there are 200 turns on the secondary.

iii. power output of the transformer.


b. A metallic conductor of length 0.40m carrying a current of 5 A experiences a 

Force of 26N in a magnetic field of strength 15 T. Determine the angle between 

the current in the conductor and the magnetic field. 


1
Expert's answer
2021-11-18T10:16:58-0500

a)

1)

"k=\frac{U_1}{U_2}=\frac{n_1}{n_2},\implies"

"n_1=n_2\frac{U_1}{U_2}=3833,"

2)

"P_2=U_2I_2=48~A,"

b)

"F=IBl\sin\alpha,\implies"

"\alpha=\arcsin \frac F{IBl}=60°."